The function 𝑓(π‘₯) is defined as 𝑓(π‘₯)=π‘Ž+𝑏cosπ‘₯, where π‘Ž and 𝑏 are constants. The range of 𝑓(π‘₯) is given by βˆ’6≀𝑓(π‘₯)≀2.

a. Find the values of π‘Ž and 𝑏

b. Solve the equation 𝑓(π‘₯)=0 for 0°≀π‘₯≀360Β°

c. Sketch the graph of 𝑦=𝑓(π‘₯) for 0°≀π‘₯≀360Β°


1
Expert's answer
2022-03-17T05:16:52-0400

a)

Because the min value of the function = -6: "a+bcosx =-6"

max value = 2: "a+bcosx = 2"

Because "-1\u2264cosx\u22641" it means "min(cosx)=-1, max(cosx)=1"

Make up a system:"\begin{cases}
 a+b(-1)=-6 \\a+b*1=2
 
\end{cases}" "\implies" "\begin{cases}
 a-b=-6 \\
 a+b=2
\end{cases}" "\implies" "\begin{cases}
 2a=-4 \\
 a+b = 2
\end{cases}" "\implies\begin{cases}
 a =-2 \\
 b=4
\end{cases}"

Answer:"f(x)=-2+4cosx"


b)

"-2+4cosx=0, x\isin[0;2\pi]"

"4cosx=2\implies cosx=\frac24 \implies cosx = \frac 12"


"x=\begin{smallmatrix}+\\-
\end{smallmatrix}" "\frac\pi3+2\pi n, n\isin Z, Z =" set of integer


for "n=0:" "x =\frac\pi3" and "x=-\frac\pi3" - not suitable

for "n =1:x=\frac\pi3+2\pi" - not suitable, and "x=-\frac\pi3 +2\pi = \frac{5\pi}3"

for "n=-1: x= \frac\pi3-2\pi" - not suitable, and "x=-\frac\pi3-2\pi" - not suitable


Answer: "x=\frac\pi3, x=\frac{5\pi}3"


c)

The required plot of the graph between the blue and green line in the screenshot